# Counter-example to the strong inequality at initial low temperature

### v4

@@ Abstract
When a thermally isolated system performs a driving process in the quasistatic regime, its variation of average work $\langle W\rangle$ is equal to its quasistatic work $\langle W_{\rm qs}\rangle$. Even though presenting this simple definition, few attempts have been made to describe such equilibrium quantity from the fluctuation theorem point of view. In this work, based on Jarzynski’s equality, four forms of such equality are deduced. To corroborate the result, a relation with the strong inequality $\langle W\rangle\geq\langle W_{\rm qs}\rangle$ is pursued. It is concluded that any of the fluctuation theorems deduced cannot derive such a postulate. Also, no contradiction is observed if the strong inequality is broken down. Based then on violations for systems starting in the microcanonical ensemble, a counter-example to the strong inequality at initial low temperature is presented.
@@ I Introduction
In this work, I explore such a problem by proving four fluctuation theorems related to the quasistatic work. All of them are deduced based on Jarzynski’s equality. However, none can derive the strong inequality. Also, there is no contradiction with previous laws if such a violation exists. Based on Ref. [22], where the strong inequality breaks down for systems starting in the microcanonical ensemble, a counter-example at initial low temperature is then presented.
@@ II Problem and proposed solution
However, four modifications in this fluctuation theorem to make it valid can be considered: 1) $\langle W_{\rm qs}\rangle$ is a type of $\Delta F$, but referring to another system than the original one; 2) the exponential is changed by another increasing and convex function; 3) the initial probability distribution is different than the canonical one; 4) Jarzynski’s equality is considered for two averages relating the work random variable $W$ and quasistatic work random variable $W_{\rm qs}$. Thus, by using Jensen’s inequality, one recovers relations where the strong inequality can be explored. That is our objective in the next sections.
@@ VI Fourth fluctuation theorem
VII Counter-example
Jarzynski’s fluctuation theorem [27] and the ones derived here can not prove the strong inequality generically. Two possible paths one faces here: the demonstration is not found out yet, or there is a counter-example where the strong inequality breaks down. In the next paragraphs, I present an example of where this second case happens.
In Ref. [29], it was shown that the quartic oscillator described by the following linear-driven Hamiltonian
$\mathcal{H}(\lambda(t))=\frac{p^{2}}{2}+\lambda(t)\frac{q^{4}}{2},\quad\lambda(t)=\lambda_{0}+\frac{t}{\tau}\delta\lambda,$ (28)
has intervals of switching times $\tau$ where the excess work $\langle W\rangle-\langle W_{\rm qs}\rangle$ is negative. The regime treated there was weakly driven perturbations, with $|t\delta\lambda/\lambda_{0}\tau|\ll 1$ for $t\in[0,\tau]$, starting with the system in a state of the microcanonical ensemble. First, our situation differs from the presented example since we consider systems starting in the canonical ensemble. Also, since the system is not in the thermodynamic limit, the equivalence between ensembles does not hold. However, when the system starts in $T=0$, the equivalence property exceptionally holds. It is expected then that at low temperatures, $T\approx 0$, the excess work calculated with the canonical ensemble will still present negative values for some switching times since Boltzmann’s factor presents a continuous dependence on the temperature. The task then becomes to find by inspection the values that lead to this breakdown.
Consider then $\beta=10^{3}$, $\lambda_{0}=1$, $\delta\lambda=0.1\lambda_{0}$, $\tau=10\lambda_{0}^{-1/2}$, and $\tau_{\rm qs}=10^{3}\lambda_{0}^{-1/2}$. Here, $\tau_{\rm qs}$ is the switching time used to calculate the quasistatic work. The procedure was the following: for 50 times, I took the average of $10^{4}$ values of $W-W_{\rm qs}$, which were numerically sampled from the initial canonical ensemble. Taking the average and standard deviation over the average values of the experiments, I obtained
$\langle W\rangle-\langle W_{\rm qs}\rangle=(-8.4\pm 0.3)\times 10^{-4}\beta^{-1},$ (29)
which shows a negative value inside the error range. Therefore, the strong inequality is broken down and the results obtained in Jarzynski’s work [27] and this one make sense. Finally, this result is exact, and no approximations of linear response theory have been used.
VIII Conclusion
I presented four fluctuation theorems related to the quasistatic work of a thermally isolated system started in a thermal state. All of them were deduced based on Jarzynski’s equality. Even though, none of them derived the strong inequality. Also, assuming the negativeness of such a difference does not imply any contradiction with already known laws and those deduced in these fluctuation theorems. Based then on violations of the strong inequality for systems starting in the microcanonical ensemble, a counter-example at initial low temperature is presented. This result corroborates the possible violations of the strong inequality obtained in the fluctuation theorems deduced in this work.

### v5

# Fluctuation theorems for quasistatic work

@@ Abstract
When a thermally isolated system performs a driving process in the quasistatic regime, its variation of average energy is equal to its quasistatic work. Even though presenting this simple definition, few attempts have been made to describe such quantity from the fluctuation theorem point of view. In this work, based on Jarzynski’s equality, four forms of such equality are deduced. To corroborate the result, a relation with the strong inequality $\langle W\rangle\geq\langle W_{\rm qs}\rangle$ is pursued. It is concluded in the end that any of the fluctuation theorems deduced cannot derive such a postulate. Also, no contradiction is observed if the strong inequality breaks down.
@@ I Introduction
In this work, I explore such a problem by proving four fluctuation theorems related to the quasistatic work. All of them are deduced based on Jarzynski’s equality. However, none can derive the strong inequality.
@@ II Problem and proposed solution
However, four modifications in this fluctuation theorem to make it valid can be considered: 1) $\langle W_{\rm qs}\rangle$ is a type of $\Delta F$, but referring to another system than the original one; 2) the exponential is changed by another increasing and convex function; 3) the initial probability distribution is different than the canonical one; 4) Jarzynski’s equality is considered for two averages relating the work random variable $W$ and quasistatic work random variable $W_{\rm qs}$. Thus, by using Jensen’s inequality, one recovers relations where the strong inequality can be explored.
@@ VI Fourth fluctuation theorem
VII Conclusion
I presented four fluctuation theorems related to the quasistatic work of a thermally isolated system. All of them were deduced based on Jarzynski’s equality. Even though, none of them derived the strong inequality. Indeed, when Jarzynski’s equality was required in these modifications, it did not offer a tight lower bound to the difference $\langle W\rangle-\langle W_{\rm qs}\rangle$. Also, assuming the negativeness of such a difference does not imply any contradiction with already known laws and those deduced in these fluctuation theorems. A piece of extra information must be provided to prove such postulate.
Conflict of interest statement
There is no conflict of interest in the parts involved.
Data availability
No data was generated in the present study.
There is no conflict of interest in the parts involved.